OpenStudy (anonymous):

(2x^3-x^2-3x)/(x^2+x) a. 2x-3x b. 2^2-3x c. 2x^2+x d. 2x-3

OpenStudy (anonymous):

\[\frac{2x^3-x^2-3x}{x^2+x}=\frac{x(2x^2-x-3)}{x(x+1)}=\frac{x(2x-3)(x+1)}{x(x+1)}=\cdots \] (eliminate the common factors in the numerator and denominator)